how do you multiply (4^-2 * 2^-3)
?

Answer by MathPro(14)

Good Afternoon,
Your problem is:
how do you multiply (4^-2 * 2^-3)?
First let’s rewrite this and use 4 = 2*2 = 2^2 to get
(2*2)^-2 * 2^-3 = (2^2)^-2*2^-3
Another rule is (x^a)^b = x ^(a*b). Lets use this rule to rewrite the equation as:
2^(-2*2) * 2^-3 =
2 ^-4 *2^-3
Use the product rule : x^a*x^b = x^(a + b) to get
2^( -4 +-3) = 2 ^ -7
Check the answer:
4^-2*2^-3 = (1/16)*(1/8)=1/128
2^-7 = 1/128
Both agree so the answer is correct.
Good Luck!
Selma from Rent-A-Whiz, an online tutoring and homework help website. Just visit www.rentawhiz.com to learn more. Currently, we are offering to solve one math problem for free. Just visit our website, select the homework tab, then click on the math problem submission link and quote coupon code 003 when you submit your problem.